# Can Performant LLMs Be Ethical? Quantifying the Impact of Web Crawling Opt-Outs

### v1

@@ Abstract
The increasing adoption of web crawling opt-outs by copyright holders of online content raises critical questions about the impact of data compliance on large language model (LLM) performance. However, little is known about how these restrictions (and the resultant filtering of pretraining datasets) affect the capabilities of models trained using these corpora. In this work, we conceptualize this effect as the data compliance gap (DCG), which quantifies the performance difference between models trained on datasets that comply with web crawling opt-outs, and those that do not. We measure the data compliance gap in two settings: pretraining models from scratch and continual pretraining from existing compliant models (simulating a setting where copyrighted data could be integrated later in pretraining). Our experiments with 1.5B models show that, as of January 2025, compliance with web data opt-outs does not degrade general knowledge acquisition (close to 0% DCG). However, in specialized domains such as biomedical research, excluding major publishers leads to performance declines. These findings suggest that while general-purpose LLMs can be trained to perform equally well using fully open data, performance in specialized domains may benefit from access to high-quality copyrighted sources later in training. Our study provides empirical insights into the long-debated trade-off between data compliance and downstream model performance, informing future discussions on AI training practices and policy decisions.
$\dagger$ denotes equal supervision
@@ 1 Introduction
One of the most notable cases in this debate is the lawsuit filed by The New York Times against OpenAI and Microsoft111nytco-assets.nytimes.com/2023/12/Lawsuit-Document-dkt-1-68-Ex-J.pdf, alleging the unauthorized use of its articles to train GPT models. The accusation is two fold: first, ChatGPT can generate verbatim excerpts from its copyrighted articles; second, ChatGPT has also hallucinated articles attributed to the Times. This case brings copyright infringement and AI development to the forefront. Amid ongoing legal and ethical debates, AI companies offer an opt-out mechanism for content publishers. Publishers can modify their Robots Exclusion Protocol (REP), commonly known as robots.txt, to block web crawlers at their discretion.
@@ 2 Related Works
Data Valuation. In the Machine Learning field, data valuation is to understand the impact of data in downstream tasks. Koh & Liang (2020) popularized the Leave-One-Out (LOO) valuation (Weisberg & Cook, 1982) from statistics to measure data impact for black-box ML approaches. Inspired by cooperative game theory, Data Shapley (Ghorbani & Zou, 2019) is another popular approach for measuring data importance, which offers a more theoretically fair attribution of data value compared to LOO. However, it is significantly more computationally intensive than LOO, as it requires retraining over all possible data subsets. While LOO offers a more tractable alternative, it still involves model retraining, making it costly for large-scale models. In the LLM world, more computationally feasible methods such as influence functions (Kwon et al., 2024; Choe et al., 2024; Xia et al., 2024) are more often used. Yet these methods often rely on gradient alignment and model loss is not directly indicative for downstream performance for LLMs. Thus, these methods can be inaccurate. In this work, aiming for a more accurate measurement, we adopt LOO for evaluating the impact of robots.txt excluded data, and we propose more computationally feasible alternatives specifically designed for LLM pretraining.
@@ 3 Data Inspection
We analyze the FineWeb-Edu dataset (Lozhkov et al., 2024), excluding any documents that, as of January 2025, disallow crawling to the bots listed in Appendix B in their robots.txt file. While FineWeb-Edu is derived from CommonCrawl444commoncrawl.org—whose data dates back as far as 2008 and respects crawling opt-outs—these opt-outs are enforced only at the time of crawling. Since data owners can update their robots.txt file at any time, CommonCrawl may contain content from websites that block crawlers at the present day but did not block at the time of crawling. Therefore, we ensure full compliance by filtering based on the most recent version and our final dataset is robots.txt-compliant. We refer to the original, unfiltered dataset as non-compliant, and the filtered version as compliant. In comparison to the non-compliant corpus, the compliant version contains 8% fewer training tokens. This ratio is consistent with what was reported in Longpre et al. (2024).
@@
We collected historical snapshots of the robots.txt files for the Top 20 domains and plotted the first recorded instance of each domain blocking a bot (as listed in Appendix B) in Figure 1. Notably, Bytespider bot by ByteDance was already being blocked as early as 2021. Since mid-2023, the OpenAI bot began to appear in blocklists. By mid 2024, more major tech companies had started facing blocks across these top domains.
@@
To assess whether adhering to the opt-out process affects the data distribution, we used WebOrganizer (Wettig et al., 2025) to cluster documents based on both topic and format, two orthogonal features representing the subject of content (topic) and the form of the content (format). We then visualized the distribution of these clusters in Figure 3. Interestingly, the overall distribution differences between compliant and non-compliant data are minimal, with the notable exception of a decrease in the proportion of News Articles and Science & Tech. and an increase in Personal Blogs.
@@
To further validate our data processing pipeline, we searched for matching 50-grams of data from nytimes.com, medicalxpress.com, and stackexchange.com domains between the compliant and non-compliant corpora. While the frequency of copyrighted articles is reduced in the compliant corpus, it is not entirely eliminated. Our manual inspection of the compliant corpus found that many remaining overlaps in nytimes.com data result from republished content, while matched medicalxpress.com content in the compliant dataset is itself republished from other sources; in contrast, stackexchange.com content primarily consists of references to other websites and is not republished. We provide examples of matched 50-grams found in the compliant dataset in Appendix C.
@@
Table 1: Percetange of documents that have a 50-gram match with documents coming from nytimes.com, medicalxpress.com, or stackexchange.com domains in compliant and non-compliant FineWeb-Edu training corpora
@@ 5.2 Compliance Gap Measured from Training from Scratch
-News 35.1 69.1 20.6 25.4 31.8 70.6 40.5 42.0 37.2 55.2 42.6
Table 2 shows no significant difference in general knowledge performance between compliant and non-compliant pre-trained models. However, Figure 5 reveals a noticeable compliance gap in structural and adversarial knowledge. This likely stems from the fact that structural formats are prevalent in documents from Math & Science and Biomedical domains, and resistance to falsified facts can be enhanced through the inclusion of newspaper articles—areas often associated with organizations that restrict web crawling.
@@
Since obtaining a comprehensive list of all URLs within a specific domain is challenging, we rely on existing sources. To simulate a scenario where all newspaper publishers opt out, we use a list of 1,158 news publishers curated by homepages.news [^5] https://palewi.re/docs/news-homepages/openai-gptbot-robotstxt.html, which serves as a reasonable proxy for the broader set of news domains. We then exclude all newspaper articles from these domains in our compliant corpus, resulting in an approximate 4% reduction in total tokens.
Interestingly, even after excluding all News domains, we observe almost no change in downstream performance, as shown in Table 2. Moreover, in the Pinocchio-Temporal and Pinocchio-Real World tasks—where newspaper articles would be expected to perform well—we observe no drop in performance even after removing all newspaper content. We believe this is due to content republishing and factual knowledge being present from other sources, as analyzed in Section 3.
Takeaway 1.
Pre-training on fully open data does not significantly impact general knowledge understanding. Even if all news publishers opt out, the effect remains minimal.
Takeaway 2.
Compared to training on non-compliant data, compliant pretraining tends to reduce a model’s ability to understand structural knowledge (i.e., information presented in structured formats) and its robustness against falsified knowledge (strategically crafted to mislead advanced LLMs).
Takeaway 3.
Newspaper articles may not be as important as expected. Temporal and real-world knowledge, which are prominently featured in newspaper articles, can be supplemented through republication and other sources.
@@ 5.3 Compliance Gap Measured from Continual Pre-Training
Major Non-compliant Domains. Based on the taxonomy in Section 3, we focus on the major domains identified among the Top 50 filtered URLs: News, Medical, and Math & Science. We categorize the Top 50 URL domains into the aforementioned three groups, and construct three domain-specific sets of non-compliant data, detailed in Appendix D.1. To perform a fine-grained ablation, we compare the performance of compliant continual pretraining with that of continual pretraining on the same compliant dataset (4.5T tokens) augmented by reintegrating domain-specific non-compliant documents. The proportion of non-compliant tokens added reflects their natural distribution in the original corpus.
@@
Table 3 presents the domain-specific cooldown results. Variations within each column excluding the underlined values can be interpreted as noise or random fluctuations. Among the three domains, only the Medical tasks show a consistent signal. A similar pattern emerges with the Pinocchio benchmarks, where reintroducing non-compliant Medical data leads to the most pronounced compliance gap, as shown in Figure 6.
@@ 1.7T Tokens Trained
Takeaway 4.
A noticeable compliance gap exists when non-compliant medical domain data are introduced at a later stage of training.
@@ 5.4 Reduced Memorization of Copyrighted Articles?
As data without crawling consent are removed from the corpus, we would expect compliant models to have less memorization of copyrighted content. To test this, we filter out all New York Times articles from the non-compliant corpus and measure the memorization of these articles across our trained models. We sampled 12’800 New York Times articles and for each article, we extracted the first 128, 256, and 512 tokens as prompts, then compared the models’ generations against the original text. Following the setup of Freeman et al. (2024), we also measure the longest common continuous substrings (LCCS), as an indication of verbatim memorization. On top of that, we add the BLEU metric, which is calculated based on N-gram overlap.
We further curated two benchmarks from these 12’800 articles, each containing 1000 questions: (1) NYTimes Multiple Choice Completion (NYTimes-MCC) task, to test if a model can select the correct completion among multiple choices given a partial excerpt from a news article. (2) NYTimes Multiple Choice Question (NYTimes-MCQ) task, which asks factual questions derived from these articles, aiming to provide challenging questions grounded in unique event details or context found in the news coverage. The benchmarks are verified using Proprietary LLMs. We offer more details regarding the scripts used for generating the QAs and more experimental results in Appendix E.
Table 4 presents the results of the memorization assessment. For comparison, we include models from the Qwen2.5 (Yang et al., 2024) and Llama3.2 (Grattafiori et al., 2024) families, noting that larger models tend to exhibit higher levels of memorization. Compared to the non-compliant model, the compliant model demonstrates reduced memorization of copyrighted content, as reflected by lower LCCS, BLEU, and NYT-MCC scores. Yet, being compliant also results in a loss of specific knowledge derived from these New York Times articles.
@@
PT + CD (Compliant with News Removed, 100B) 21.06 0.55 48.9 27.3
@@
Takeaway 5.
Compliant training lowers verbatim memorization of copyrighted content but also results in a loss of specific knowledge derived from the publications.
@@ 6 Conclusions and Future Work
In this work, we take an initial step toward examining the widely discussed trade-off between data compliance and downstream model performance. We introduce the concept of the Data Compliance Gap (DCG) to quantify performance differences resulting from adhering to robots.txt opt-outs, and evaluate DCG across various training settings. Our findings indicate that it is currently feasible to train performant LLMs exclusively on data that respect web crawling restrictions. Thus, we recommend adherence to robots.txt restrictions for LLM developments.
@@ Acknowledgement
DF would like to thank Maksym Andriushchenko for his feedback on an initial draft of the paper. AB gratefully acknowledges the support of the Swiss National Science Foundation (No. 215390), Innosuisse (PFFS-21-29), the EPFL Center for Imaging, Sony Group Corporation, and a Meta LLM Evaluation Research Grant. This work was supported as part of the Swiss AI Initiative by a grant from the Swiss National Supercomputing Centre (CSCS) under project ID a06 on Alps.
@@ Appendix B List of blocked web crawlers
Appendix C Matched 50-grams in the compliant dataset

### v2

@@ Abstract
The increasing adoption of web crawling opt-outs by copyright holders of online content raises critical questions about the impact of data compliance on large language model (LLM) performance. However, little is known about how these restrictions (and the resultant filtering of pretraining datasets) affect the capabilities of models trained using these corpora. In this work, we conceptualize this effect as the data compliance gap (DCG), which quantifies the performance difference between models trained on datasets that comply with web crawling opt-outs, and those that do not. We measure the data compliance gap in two settings: pretraining models from scratch and continual pretraining from existing compliant models (simulating a setting where copyrighted data could be integrated later in pretraining). Our experiments with 1.5B models show that, as of January 2025, compliance with web data opt-outs does not degrade general knowledge acquisition (close to 0% DCG). However, in specialized domains such as biomedical research, excluding major publishers leads to performance declines. These findings suggest that while general-purpose LLMs can be trained to perform equally well using fully open data, performance in specialized domains may benefit from access to high-quality copyrighted sources later in training. Our study provides empirical insights into the long-debated trade-off between data compliance and downstream model performance, informing future discussions on AI training practices and policy decisions. Our website is available at https://data-compliance.github.io/. $\dagger$ denotes equal supervision
@@ 1 Introduction
One of the most notable cases in this debate is the lawsuit filed by The New York Times against OpenAI and Microsoft111nytco-assets.nytimes.com/2023/12/Lawsuit-Document-dkt-1-68-Ex-J.pdf, alleging the unauthorized use of its articles to train GPT models. The accusation is two-fold: first, ChatGPT can generate verbatim excerpts from its copyrighted articles; second, ChatGPT has also hallucinated articles attributed to the Times. This case brings copyright infringement and AI development to the forefront. Amid ongoing legal and ethical debates, AI companies offer an opt-out mechanism for content publishers. Publishers can modify their Robots Exclusion Protocol (REP), commonly known as robots.txt, to block web crawlers at their discretion.
@@ 2 Related Works
Data Valuation. In the Machine Learning field, data valuation is to understand the impact of data in downstream tasks. Koh & Liang (2020) popularized the Leave-One-Out (LOO) valuation (Weisberg & Cook, 1982) from statistics to measure data impact for black-box ML approaches. Inspired by cooperative game theory, Data Shapley (Ghorbani & Zou, 2019) is another popular approach for measuring data importance, which offers a more theoretically fair attribution of data value compared to LOO. However, it is significantly more computationally intensive than LOO, as it requires retraining over all possible data subsets. While LOO offers a more tractable alternative, it still involves model retraining, making it costly for large-scale models (Fan et al., 2025). In the LLM world, more computationally feasible methods such as influence functions (Kwon et al., 2024; Choe et al., 2024; Xia et al., 2024) are more often used. Yet these methods often rely on gradient alignment and model loss is not directly indicative for downstream performance for LLMs. Thus, these methods can be inaccurate. In this work, aiming for a more accurate measurement, we adopt LOO for evaluating the impact of robots.txt excluded data, and we propose more computationally feasible alternatives specifically designed for LLM pretraining.
@@ 3 Data Inspection
We analyze the FineWeb-Edu dataset (Lozhkov et al., 2024), excluding any documents that, as of January 2025, disallow crawling to the bots listed in Appendix B in their robots.txt file. While FineWeb-Edu is derived from CommonCrawl444commoncrawl.org—which dates back to 2008 and respects robots.txt restrictions—those restrictions are enforced only at the time of crawling. Since website owners can modify their robots.txt files at any time, CommonCrawl may include content from sites that currently block crawlers but did not at the time of data collection. To ensure full compliance, we apply retrospective filtering using the most recent robots.txt files, resulting in a dataset that adheres to current crawling restrictions. We refer to the original, unfiltered dataset as non-compliant, and the filtered version as compliant. In comparison to the non-compliant corpus, the compliant version contains 8% fewer training tokens. This ratio is consistent with what was reported in Longpre et al. (2024).
@@
We collected historical snapshots of the robots.txt files for the Top 20 domains and plotted the first recorded instance of each domain blocking any of the AI training crawling bots (as listed in Appendix B) in Figure 1. Notably, Bytespider bot by ByteDance was already being blocked as early as 2021. Since mid-2023, the OpenAI bot began to appear in blocklists. By mid 2024, more major tech companies had started facing blocks across these top domains.
@@
To assess whether adhering to the opt-out process affects the data distribution, we used WebOrganizer (Wettig et al., 2025) to cluster documents based on both topic and format, two orthogonal features representing the subject of content (topic) and the form of the content (format). We then visualized the distribution of these clusters in Figure 3. Interestingly, the overall distribution differences between compliant and non-compliant data are minimal, with the notable exception of a decrease in the proportion of News Articles and Science & Tech. and an increase in Personal Blogs. For completeness, the percentage decrease in each category is presented in Appendix C.
@@
To further validate our data processing pipeline, we searched for matching 50-grams of data from nytimes.com, medicalxpress.com, and stackexchange.com domains between the compliant and non-compliant corpora. While the frequency of copyrighted articles is reduced in the compliant corpus, it is not entirely eliminated. Our manual inspection of the compliant corpus found that many remaining overlaps in nytimes.com data result from republished content, while matched medicalxpress.com content in the compliant dataset is itself republished from other sources; in contrast, stackexchange.com content primarily consists of references to other websites and is not republished. We provide examples of matched 50-grams found in the compliant dataset in Appendix D.
Table 1: Percentage of documents that have a 50-gram match with documents coming from nytimes.com, medicalxpress.com, or stackexchange.com domains in compliant and non-compliant FineWeb-Edu training corpora
@@ 5.2 Compliance Gap Measured from Training from Scratch
[…]
Table 5.2 shows no significant difference in general knowledge performance between compliant and non-compliant pre-trained models. However, Figure 5 reveals a noticeable compliance gap in structural and adversarial knowledge. This likely stems from the fact that structural formats are prevalent in documents from Math & Science and Biomedical domains, and resistance to falsified facts can be enhanced through the inclusion of newspaper articles and scientific papers—areas often associated with organizations that restrict web crawling.
@@
Since obtaining a comprehensive list of all URLs within a specific domain is challenging, we rely on existing sources. Currently, a critical domain is newspapers. To simulate a scenario where all newspaper publishers opt out, we use a list of 1,158 news publishers curated by homepages.news [^5] https://palewi.re/docs/news-homepages/openai-gptbot-robotstxt.html, which serves as a reasonable proxy for the broader set of news domains. We then exclude all newspaper articles from these domains in our compliant corpus, resulting in an approximate 4% reduction in total tokens. As shown in Appendix D, the republication of news articles is quite common. To accurately assess their impact, we further perform a decontamination step that removes document segments replicated from news articles hosted on different URL domains. This process results in a further 14% reduction in total compliant tokens. After this process, the proportion of documents containing a 50-gram match with content from nytimes.com drops to 0.012%, indicating the effectiveness of the decontamination.
Interestingly, even after excluding all newspaper articles, we observe almost no change in downstream performance, as shown in Table 5.2. Moreover, in the Pinocchio-Temporal and Pinocchio-Real World tasks—where newspaper articles would be expected to perform well—we observe no drop in performance even after removing all newspaper content. We believe this is due to factual knowledge being present from other sources.
Takeaway 1. Pre-training on fully open data does not significantly impact general knowledge understanding. Even if all news publishers opt out, the effect remains minimal.
Takeaway 2. Compared to training on non-compliant data, compliant pretraining tends to reduce a model’s ability to understand structural knowledge (i.e., information presented in structured formats) and its robustness against falsified knowledge (strategically crafted to mislead advanced LLMs).
Takeaway 3. Newspaper articles may not be as important as expected. Temporal and real-world knowledge, which are prominently featured in newspaper articles, can be supplemented through other sources.
@@ 5.3 Compliance Gap Measured from Continual Pre-Training
Major Non-compliant Domains. Based on the taxonomy in Section 3, we focus on the major domains identified among the Top 50 filtered URLs: News, Medical, and Math & Science. We categorize the Top 50 URL domains into the aforementioned three groups, and construct three domain-specific sets of non-compliant data, detailed in Appendix E.1. To perform a fine-grained ablation, we compare the performance of compliant continual pretraining with that of continual pretraining on the same compliant dataset (4.5T tokens) augmented by reintegrating domain-specific non-compliant documents. The proportion of non-compliant tokens added reflects their natural distribution in the original corpus.
@@
Table 3 presents the domain-specific cooldown results. The underlined scores correspond to runs where task-specific, non-compliant data was added during cooldown, which is supposed to enhance performance on that specific task. The remaining entries illustrate expected variability across cooldown runs without such targeted augmentation. Among the three domains, only the Medical tasks with long training (1.7T tokens) show a consistent signal. A similar pattern emerges with the Pinocchio benchmarks, where reintroducing non-compliant Medical data leads to the most pronounced compliance gap, as shown in Figure 6.
[…]
@@ 1.7T Tokens Trained
Takeaway 4. A noticeable compliance gap exists when non-compliant medical domain data is introduced at a later stage of training.
@@ 5.4 Reduced Memorization of Copyrighted Articles?
As data without crawling consent is removed from the corpus, we would expect compliant models to have less memorization of copyrighted content. To test this, we filter out all New York Times articles from the non-compliant corpus and measure the memorization of these articles across our trained models. We sampled 12’800 New York Times articles and for each article, we extracted the first 128, 256, and 512 tokens as prompts, then compared the models’ generations against the original text. Following the setup of Freeman et al. (2024), we also measure the longest common continuous substrings (LCCS), as an indication of verbatim memorization. On top of that, we add the BLEU metric, which is calculated based on N-gram overlap.
We further curated two benchmarks from these 12’800 articles, each containing 1000 questions: (1) NYTimes Multiple Choice Completion (NYTimes-MCC) task, to test if a model can select the correct completion among multiple choices given a partial excerpt from a news article. (2) NYTimes Multiple Choice Question (NYTimes-MCQ) task, which asks factual questions derived from these articles, aiming to provide challenging questions grounded in unique event details or context found in the news coverage. The benchmarks are verified using Proprietary LLMs. We offer more details regarding the scripts used for generating the QAs and more experimental results in Appendix F.
Table 4 presents the results of the memorization assessment. For comparison, we include models from the Qwen2.5 (Yang et al., 2024) and Llama3.2 (Grattafiori et al., 2024) families, noting that larger models tend to exhibit higher levels of memorization. The compliant model demonstrates reduced memorization of copyrighted content compared to the non-compliant model, as shown by lower LCCS, BLEU, and NYT-MCC scores—a trend consistent with the removal of newspaper articles. Importantly, compliance does not necessarily lead to a loss of specific knowledge from New York Times articles, as indicated by comparable NYT-MCQ scores. Notably, even with news decontamination applied, the NYT-MCQ score does not decrease.
@@
PT + CD (Compl., with News Removed, 100B) 20.28 0.53 48.0 28.6
@@
Takeaway 5. Compliant training reduces verbatim memorization of copyrighted content without necessarily compromising specific knowledge derived from the publications.
@@ 6 Conclusions and Future Work
In this work, we take an initial step toward examining the widely discussed trade-off between data compliance and downstream model performance. We introduce the concept of the Data Compliance Gap (DCG) to quantify performance differences resulting from adhering to robots.txt opt-outs, and evaluate DCG across various training settings. Our findings indicate that it is currently feasible to train performant LLMs exclusively on data that respects web crawling restrictions. Thus, we recommend adherence to robots.txt restrictions for LLM developments.
We clarify that our notion of compliance addresses only part of data usage rights. Our approach focuses on AI-specific user agents, offering a clearer signal of consent than general crawling permissions. However, it does not cover all legal constraints, such as Terms of Service in “browsewrap agreements”, which are legally ambiguous and challenging to process at scale. Therefore, our robots.txt filtering represents a practical upper bound on demonstrable compliance, not full legal clearance. We call for future work for a finer-grained study.
@@ Acknowledgement
DF would like to thank Maksym Andriushchenko for his feedback on an initial draft of the paper. DF acknowledges funding from SNSF Grant number 10005248. AB gratefully acknowledges the support of the Swiss National Science Foundation (No. 215390), Innosuisse (PFFS-21-29), the EPFL Center for Imaging, Sony Group Corporation, and a Meta LLM Evaluation Research Grant. This work was supported as part of the Swiss AI Initiative by a grant from the Swiss National Supercomputing Centre (CSCS) under project ID a06 on Alps.
@@ Appendix A Limitations
We acknowledge that robots.txt compliance addresses only one aspect of data usage rights. Our approach evaluates robots.txt rules specifically for AI training user agents, offering a clearer signal of consent than general crawling permissions. However, this does not account for all legal restrictions, such as Terms of Service in “browsewrap agreements.” These terms are legally ambiguous and difficult to process at scale due to their natural language format. Thus, our robots.txt filtering reflects a practical upper bound of demonstrable compliance, not comprehensive legal clearance.
@@ Appendix B List of blocked web crawlers
Appendix C Percentage decrease in each category due to data compliance
In addition to changes in data distribution, Tables LABEL:tab:_percentage-drop-topic and LABEL:tab:_percentage-drop-format report the percentage of data removed by domain and relative to the entire corpus. For instance, filtering out $11.39\%$ of Science & Tech data leads to a $2.6483\%$ reduction in total corpus tokens.
Table 5: Topic domain removal percentages
Topic Domain % Removed (Domain) % Removed (Corpus)
Science & Tech. 11.39 2.6483
Health 9.65 1.8893
History 8.05 0.7921
Education & Jobs 7.11 0.7639
Politics 10.45 0.4778
Literature 8.94 0.3639
Industrial 8.02 0.2438
Home & Hobbies 6.01 0.2002
Finance & Business 8.46 0.1894
Art & Design 8.08 0.1832
Transportation 10.84 0.1676
Religion 3.58 0.1377
Software Dev. 6.52 0.1327
Crime & Law 8.03 0.1184
Food & Dining 8.46 0.1114
Software 6.92 0.1096
Sports & Fitness 7.36 0.0903
Hardware 9.12 0.0829
Entertainment 9.44 0.0806
Social Life 6.72 0.0572
Travel 6.94 0.0534
Games 11.69 0.0478
Fashion & Beauty 3.61 0.0089
Adult 0.81 0.0001
Table 6: Format domain removal percentages
Format Domain % Removed (Domain) % Removed (Corpus)
News Article 24.09 2.4956
Knowledge Article 6.87 1.6224
Nonfiction Writing 9.76 1.2853
Academic Writing 9.57 0.7303
Tutorial 5.95 0.5673
Product Page 6.53 0.4566
Structured Data 17.39 0.3920
Q&A Forum 27.33 0.3518
Listicle 6.62 0.2115
Content Listing 8.51 0.1708
Personal Blog 1.68 0.1330
Truncated 10.22 0.1236
News (Org.) 2.49 0.0942
Audio Transcript 12.23 0.0940
Documentation 7.58 0.0804
FAQ 5.13 0.0336
Comment Section 4.13 0.0233
Customer Support 2.88 0.0223
About (Org.) 1.15 0.0208
Creative Writing 4.09 0.0154
Spam / Ads 2.17 0.0122
User Review 4.09 0.0103
About (Pers.) 1.85 0.0023
Legal Notices 1.67 0.0020
Appendix D Matched 50-grams in the compliant dataset